### v2

# A Novel Quantum Realization of Jet Clustering in High-Energy Physics Experiments

A Novel Quantum Realization of Jet Clustering in High-Energy Physics Experiments
Yongfeng Zhu WeiFeng Zhuang Chen Qian Yunheng Ma Dong E. Liu dongeliu@mail.tsinghua.edu.cn Manqi Ruan ruanmq@ihep.ac.cn Chen Zhou czhouphy@pku.edu.cn
journal: Science Bulletin
[…]
[inst1]organization=State Key Laboratory of Nuclear Physics and Technology, School of Physics, Peking University, addressline=Haidian District, city=Beijing, postcode=100871, country=China
[…]
[inst2]organization=Beijing Academy of Quantum Information Sciences, addressline=Haidian District, city=Beijing, postcode=100193, country=China
[…]
[inst3]organization=Department of Physics, Tsinghua University,addressline=Haidian District, city=Beijing, postcode=100084, country=China
[…]
[inst4]organization=Institute of High Energy Physics, Chinese Academy of Sciences, addressline=Shijingshan District, city=Beijing, postcode=100049, country=China
[…]
[inst5]organization=University of Chinese Academy of Sciences, addressline=Shijingshan District, city=Beijing, postcode=100049, country=China
Exploring the application of quantum technologies to fundamental sciences holds the key to fostering innovation for both sides. In the near future, the number of reliable quantum operations will be constrained by noise and decoherence [1], which refers to any disturbance causing errors in the quantum state of qubits. To overcome these challenges, hybrid quantum-classical algorithms [2] have been proposed. A notable example is the Quantum Approximate Optimization Algorithm (QAOA) [3], an algorithm commonly used to address classical combinatorial optimization problems. In high-energy particle collisions, quarks and gluons are produced and immediately form collimated particle sprays known as jets due to color confinement [4]. Accurate jet clustering is crucial as it retains the originating quark or gluon information and forms the basis for studying properties of the Higgs boson, which underlies the mechanism of mass generation for subatomic particles. For the first time, we leverage QAOA for jet clustering, as shown in Fig. 1e, by mapping collision events into graphs, as shown in Fig. 1d. We obtain experimental results on quantum computer simulators and quantum computer hardware, whose performances are compared to or even better than classical algorithms for a small-sized problem.
Figure 1: A typical physics analysis procedure involving jets in high-energy physics is outlined as follows: Panel (a) [5] depicts the layout of the Circular Electron-Positron Collider. Electrons and positrons are accelerated in the booster and collide at the Interaction Point (IP) within the detector, as illustrated in panel (b) [5]. Resultant quarks and gluons from these collisions transform into collimated streams of particles known as jets. Panel (c) illustrates an event $e^{+}e^{-}\to ZH(Z\to\nu\bar{\nu},H\to s\bar{s})$, where two quarks form two jets. Jets are clustered using a jet clustering algorithm, a method that groups final-state particles into sets corresponding to individual gluons or quarks. During jet clustering, each event can be represented as a graph where particles are nodes and the angles between particles are edges, as shown in panel (d). The Quantum Approximate Optimization Algorithm (QAOA) can be applied to perform jet clustering, as demonstrated in panel (e). Based on the architecture of the quantum processor and available elementary gates, the compiled quantum circuit is shown in panel (f). The information from reconstructed jets can be utilized to measure related objectives, such as the $H\to s\bar{s}$ measurement shown in panel (g) [6].
A typical physics analysis procedure involving jets in high-energy physics is described as follows. In an electron-positron Higgs factory, such as the Circular Electron-Positron Collider (CEPC) [5], whose layout is shown in Fig. 1a, the golden process for studying Higgs properties is $e^{+}e^{-}\to ZH$. Electrons and positrons are accelerated in the booster and interact at the Interaction Point (IP), which is located within the detector, as illustrated in Fig. 1b. An $e^{+}e^{-}\to ZH(Z\to\nu\bar{\nu},H\to s\bar{s})$ event, where two quarks form two jets, is displayed in Fig. 1c. Subsequently, a jet clustering algorithm, which groups final-state particles into sets corresponding to individual gluons or quarks, is applied to identify these jets. Following jet clustering and additional techniques, various physics analyses can be performed, such as the measurement of the branching ratio of $H\to s\bar{s}$, shown in Fig. 1g. Given the continuous progress in quantum computing technology, there is increasing interest in identifying practical applications for near-term quantum machines that aim to achieve faster solutions and improve the quality of those solutions. For instance, [7] demonstrated the formulation of thrust–whose optimum corresponds to the most jet-like separating plane defining two hemisphere jets–both as a quantum annealing problem and a Grover search problem. Additionally, other studies [8, 9] have developed quantum annealing-based jet clustering algorithms.
The QAOA draws inspiration from the Quantum Adiabatic Algorithm (QAA). QAA uses two Hamiltonians: an initial Hamiltonian $\hat{H}_{M}$ with an easily prepared ground state, and a final Hamiltonian $\hat{H}_{C}$ whose ground state encodes the solution to the optimization problem. The adiabatic evolution path is defined by the transitional Hamiltonian $\hat{H}(t)=s(t)\hat{H}_{C}+(1-s(t))\hat{H}_{M}$, where $s(t)=\frac{t}{T}$, $t\in[0,T]$, and T is the total evolution time. The evolution operator $\hat{U}(t)$ is derived from $\hat{H}(t)$. QAA can be emulated on a gate-based quantum computer by Trotterizing $\hat{U}(t)$ into sufficiently small steps,
$\begin{split}\hat{U}(t)&\approx\prod_{j=1}^{P}e^{-i\hat{H}(j\Delta t)\Delta t}=\prod_{j=1}^{P}e^{-i(1-s(j\Delta t))\hat{H}_{M}\Delta t}e^{-is(j\Delta t)\hat{H}_{C}\Delta t}\\ &=\prod_{j=1}^{P}e^{-i\beta_{j}\hat{H}_{M}\Delta t}e^{-i\gamma_{j}\hat{H}_{C}\Delta t}=\prod_{j=1}^{P}\hat{U}_{M}(\beta_{j})\hat{U}_{C}(\gamma_{j})\end{split}$ (1)
where $\Delta t=T/P$ and $P$ is the number of steps. QAOA, Inspired by this Trotterized version, was designed with repeated cost $\hat{U}_{C}(\gamma)$ and mixer layers $\hat{U}_{M}(\beta)$ [3].
An undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges, can represent a collision event in high-energy physics. Particles as nodes and the relationship between particles i and j is an edge with weight $w_{ij}$ calculated as the angle between them. For each node, edges are sorted by weight in decreasing order and only the top k edges are retained (k-regular graph). Figure 1e illustrates an $e^{+}e^{-}\to ZH(Z\to\nu\bar{\nu},H\to s\bar{s})$ event with 30 particles and k=3. After mapping collision events as graphs, QAOA can perform jet clustering akin to solving the Max-K-cut problem [10]. This study clusters two jets, analogous to the Max-cut problem (a specific case of the Max-K-cut problem with K equal to 2), aiming to partition the graph nodes into two complementary subsets to maximize the weighted sum of edges connecting nodes in different subsets, given by $C(x)=\sum_{i,j=1}^{|v|}w_{ij}x_{i}(1-x_{j})$. Maximizing $C(x)$ is NP-hard, but QAOA can approximate $C(x)$ close to its maximum value $C_{max}=\mathop{\text{max}}\limits_{x}C(x)$ [3]. To solve the Max-cut problem using QAOA, we begin by defining the problem Hamiltonian as $\hat{H}_{C}=\frac{1}{2}\sum_{(i,j)\in E}W_{ij}(I-\sigma^{z}_{i}\sigma^{z}_{j})$ and initial Hamiltonian as $\hat{H}_{M}=\sum_{j=1}^{n}\sigma_{j}^{x}$, where $\sigma^{z}_{i}$ represents the Pauli-Z operator acting on the i-th qubit and $W_{ij}$ represents the angle between particles i and j. The quantum circuit is then constructed based on the evolution operator as shown in Equation 1. Detailed QAOA procedures for the Max-cut problem are described in [3, 12]. Key aspects include the number of Trotterized steps (QAOA depth), evolution times of each step optimized by a classical optimization algorithm (COBYLA [11] and an interpolation method [12]), and sampling times (1024 in this study), as the expectation value of $C(x)$ is obtained by sampling quantum states multiple times. For practical jet clustering, the focus is on approximating the sampling state to the ground state. Among the 1024 sampling states, the one closest to the ground state is selected. Jet clustering performance is evaluated by the angle between the reconstructed jet and its corresponding quark. In the process $e^{+}e^{-}\to ZH(Z\to\nu\bar{\nu},H\to s\bar{s})$, each event produces two jets, resulting in angles $angle_{1}$ and $angle_{2}$. This study uses the cumulative angle $=angle_{1}+angle_{2}$ to assess jet clustering performance.
Figure 2: Panels (a), (b), and (c) show the jet clustering performance for 4000 events with 30 particles. (a) The jet clustering performance with k set to 6 and QAOA depths of 1, 3, and 5. (b) The performance with QAOA depth set to 3 and k values of 2, 4, 6, 7, and 8. (c) Compare the QAOA method (with QAOA depth of 3 and k of 7), $e^{+}e^{-}k_{t}$, and k-Means method. (d) Compiled quantum circuit of QAOA for performing jet clustering on a 6-particle event, based on the Baihua quantum hardware. The circuit has a depth of 26, with 34 CNOT gates and 27 single-qubit gates. (e) Jet clustering performance for 1217 6-particle events on the Quafu platform, comparing the quantum simulator, $e^{+}e^{-}k_{t}$ algorithm, and k-Means, with QAOA depth set to 1 and k set to 2.
The QAOA performance can be enhanced by increasing QAOA depth, corresponding to a more granular unitary totter evolution $\hat{U}(t)$ and allowing for finer evolutionary increments. This granularity aids in retaining the quantum system’s state within the vicinity of the ground state throughout the transition from the initial Hamiltonian to the problem Hamiltonian. To ensure statistical robustness, we analyze 4000 events comprising 30 particles. Figure 2a compares jet clustering performance at depths of 1, 3, and 5 with k = 6, where the results for depths 5 and 3 are comparable and noticeably superior to that of depth 1, confirming the successful sampling of the ground state.
A higher value of k signifies that a node is linked to more nodes within the graph, indicating a more complex complied quantum circuit. In particular, a larger k-value may lead to linking of a particle to other particles within the same jet, enhancing the intricacy of the captured relationships. Conversely, a smaller k-value may result in suboptimal performance due to inadequate linkage between particles belonging to distinct jets, thus failing to accurately cluster particles from separate jets. Figure 2b compares the performance of the jet clustering at k-values of 2, 4, 6, 7, and 8 when the depth of QAOA is 3. It shows that the jet clustering performance is better at the k-value of 7.
For a more intuitive comparison of the QAOA’s impact on jet clustering performance, Fig. 2c displays the performance of QAOA (when QAOA depth taking values of 3 and k being 7) together with two classical algorithms, k-Means and $e^{+}e^{-}k_{t}$ [13]. The k-Means is a clustering method that partitions a dataset into k clusters by minimizing the sum of squared distances between data points and the centroids of their respective clusters. For our jet clustering study, the angle between two particles is used by k-Means to describe the distance. The $e^{+}e^{-}k_{t}$ is a classical jet clustering algorithm used in electron-positron collisions, which calculates the distance between particles based on their angles and energies $d_{ij}=2mn(E_{i}^{2},E_{j}^{2})(1-cos\theta_{ij})$, where $E_{i/j}$ represents the energy of particle i/j and $\theta_{ij}$ represents the angle between particles i and j. The $e^{+}e^{-}k_{t}$ identifies the pair of closest particles recombines them, and then repeats the procedure over and over again until some stopping criterion is reached. The performance of QAOA is observed to be comparable to that of the $e^{+}e^{-}k_{t}$ algorithm and better than that of the k-Means. This comparison highlights the potential of the QAOA in the jet clustering problem.
Furthermore, we conduct the analysis on the BAQIS Quafu quantum computing cloud [14], a stable and open-access platform providing users with access to multiple superconducting quantum processors. Specifically, the Baihua processor, whose architecture can be found in [14], features 123 operational qubits interconnected through couplers. The relaxation time $T_{1}$ and the dephasing time $T_{2}^{\ast}$ typically characterize the decoherence of the qubit state. This chip has an average $T_{1}$ of 73.994 microseconds and an average $T_{2}^{\ast}$ of 29.02 microseconds [14]. Single-qubit gate operations achieve an average fidelity (probability of correct output) of 99.9% and two-qubit gates (CZ) realize an average fidelity of 98.8% [14]. Based on the architecture of the Baihua processor, such as qubit connectivity [1], and available elementary gates, the quantum circuit can be compiled with the Quafu-Qcover compiler [15] according to the quantum algorithm, specifically QAOA in this study. The compiled quantum circuit is then transpiled into a series of qubit operation commands and sent to the quantum control system of the superconducting computer. The quantum control system executes these commands sequentially, with each command corresponding to a pulse operation used to control the qubits.
An ideal quantum processor, free from noise and with sufficient coherence time, can utilize all available qubits for computations. However, due to current hardware limitations such as noise and limited connectivity–where each qubit can only interact with its immediate 2 or 3 neighbors, leading to a greater depth in compiled quantum circuits–we restrict our study to 6 qubits. For events involving 6 particles, we found that the QAOA performance is already optimal when the QAOA depth is set to 1 and k to 2. Figure 2d illustrates a compiled quantum circuit of QAOA for performing jet clustering on a 6-particle event. The compiled circuit reaches a depth of 26 with 34 CNOT gates and 27 single-qubit gates. As a demonstration, Fig. 2e showcases the performance of 1217 6-particle events conducted on the Baihua processor, illustrating that for this small-sized problem, the quantum hardware can achieve similar performance to a noiseless quantum computer simulator.
In summary, the rapid development of quantum algorithms and hardware devices enables the execution of small-scale but representative applications on quantum computers for fundamental sciences. For example, [16] introduced the QCSH package, offering an efficient quantum algorithm for calculating nuclear structure. Similarly, [17] presented a quantum algorithm for heat conduction that significantly outperforms classical counterparts, Additionally, [18] proposed the first algorithm for simulating ocean circulation on a quantum computer, achieving computational speedup, and [19] designed a quantum circuit for implementing spectral clustering with significant speedup. For the first time, by mapping collision events into graphs, we have obtained promising results in applying QAOA to jet clustering. Simulating 30 qubits on quantum computers, we found that QAOA performance, based on the angle between the reconstructed jet and the corresponding quark, is close to the $e^{+}e^{-}k_{t}$ algorithm and superior to k-Means. Furthermore, evaluating QAOA using 6 qubits on the BAQIS Quafu quantum computing cloud demonstrated that quantum hardware can achieve performance similar to a noiseless quantum simulator for a small-sized problem. This study aims to harness the potential of quantum computing technologies for reconstruction problems in high-energy physics experiments. Conversely, it provides a platform for researching and developing quantum computing technologies. Continued advancements in quantum hardware and algorithms will not only enable us to perform jet clustering on larger problems but also revolutionize our approach to complex scientific challenges, driving innovation and expanding our understanding of the universe.
Conflict of interest
The authors declare that they have no conflict of interest.
Acknowledgment
This work is supported by the Fundamental Research Funds for Central Universities, Peking University, the National Key R&D Program of China under Contracts No. 2022YFE0116900, the National Natural Science Foundation of China (NSFC) under Grant No. 12305010, and the Beijing Natural Science Foundation under Grant No. Z220002.
We thank Hongze Xu for his invaluable support in utilizing Quafu Cloud. We extend our appreciation to the Beijing Academy of Quantum Information Sciences for access to quantum hardware and to the Computing Center at the Institute of High-Energy Physics for providing computing resources.
Author contributions
Yongfeng Zhu analyzed the data and wrote the initial manuscript. Weifeng Zhuang, Chen Qian and Yunheng Ma researched QAOA and discussed it with Yongfeng Zhu. Dong E. Liu, Manqi Ruan, and Chen Zhou conceived and oversaw the project. All authors contributed to analyzing the results and to revising and reviewing the manuscript.